OpenStudy (anonymous):

What is the multiplication table for Z[ω]/(2), ω=(-1-i√3)/2?

OpenStudy (fibonaccichick666):

I'm sorry, I've never heard the term multiplication table in this sense, can you elaborate?

OpenStudy (kainui):

 Here I'll help you start, just fill in the boxes where 1 and 1 meet you write whateve 1*1= 1*ω=ω so I wrote that in too. Keep going.